OpenStudy (anonymous):

What are the first four terms of the geometric sequence with a t1 = 2 and tn = 3tn-1?

OpenStudy (anonymous):

\[ t_1=2\\ t_n= 3 t_{n-1}\\ t_2= 6\\ t_3= 3(6) =18\\ t_4= 3(18)=54\\ \]